Oneday, 45 frogs were captured from a pond, marked and then returned to the pond. The next day, 40 frogs were observed in the pond, of which 10 had been marked the previous day.Assuming the marked frogs are equally distributed among all the frogs in the pond, what is the best estimate of how many frogs live in the pond?
Answer by stanbon(75887)

Check the site http://cybersalmon.fws.gov
Click on "Activities"; your problem is listed.
Your answer would be [45*40]/10 = 1800/10= 180 fish
Cheers,
Stan H.